OpenStudy (coconutjj):

How do you input arcsec onto a calculator? is it 1/arccos ?

OpenStudy (kainui):

Almost, but not quite. I just sorta figured this out now by playing around with the algebra. So if you have \[\Large \sec \theta = x \\ \Large \theta = arcsec x\] but you can't do that cause it doesn't exist on your calculator so you have to turn it into: \[\Large \sec \theta = x \\ \Large \frac{1}{\cos \theta} = x \\ \Large \frac{1}{x} = \cos \theta \\ \Large \arccos( \frac{1}{x} ) = \theta\] so if you forget again you can just figure it out in the same way I did here, it's just 1/ on the inside of arccosine.

OpenStudy (kainui):

I guess I should have included the final result equal to theta as the same as the first result equal to theta: \[\Large \theta = arcsec x \\ \Large \theta = \arccos \left( \frac{1}{x} \right) \\ \Large arcsec( x) = \arccos ( \frac{1}{x})\]

OpenStudy (coconutjj):

I see... Thank You ! :)

OpenStudy (kainui):

Yeah now you can figure it out for others too like arccot or arccsc if you need to =P